OpenStudy (anonymous):

Classify the solutions of 1 over x plus 4, plus one half, equals 1 over x plus 4 as extraneous or non-extraneous

OpenStudy (misty1212):

lol there is either a solution, or there is not not sur what math teacher made up this "extraneous' nonsense

OpenStudy (misty1212):

\[\frac{1}{x+4}+\frac{1}{2}=\frac{1}{x+4}\]? is that the question?

OpenStudy (anonymous):

i just need help with the extraneous and non extraneous

OpenStudy (anonymous):

yes that is

OpenStudy (misty1212):

a bit hard to read

OpenStudy (misty1212):

oh ok then lets think

OpenStudy (misty1212):

x is a variable, it can be anything but whatever it is on the left, it is the same on the right so if you add one half to a number, it is impossible to get the same number back there is NO SOLUTION

OpenStudy (anonymous):

the answer is x= -4 but idk if its extraneous or not

OpenStudy (misty1212):

THE ANSWER IS NOT X = -4

OpenStudy (anonymous):

then what is it? bc i followed the directions in the lesson

OpenStudy (misty1212):

ok fine, say \(x=-4\) is "extraneous" if that will make your math teacher happy

OpenStudy (anonymous):

do you know the difference between the two?

OpenStudy (misty1212):

\[\overbrace{\frac{1}{x+4}}^{\text{this}}+\frac{1}{2}=\overbrace{\frac{1}{x+4}}^{\text{is equal to this}}\]

OpenStudy (anonymous):

ok

OpenStudy (misty1212):

you cannot add one half to a number and get the same number back

OpenStudy (misty1212):

the not very bright math teacher who wants you not to think like a person, but do some steps to "solve" wants you to find that \(x=-4\) but \(x\) cannot be \(-4\) because a) that would make the denominator zero and b) there is NO SOLUTION

OpenStudy (misty1212):

therefore they want you to behave like a robot and say "the solution is \(x=-4\) but it is "extraneous""

OpenStudy (anonymous):

x = −4; extraneous x = −4; non-extraneous x = −8; extraneous x = −8; non-extraneous so out of these answer choices......its the first one?

OpenStudy (anonymous):

@misty1212